# MEGS2: Memory-Efficient Gaussian Splatting via Spherical Gaussians and Unified Pruning

### v1

@@ 4.3 Qualitative Results
| Method | PSNR | SSIM | LPIPS | VRAM |
| GaussianSpa + Reduce3DGS | 402 | |||
| GaussianSpa (SH SG) | 339 | |||
| Ablation 1: soft hard | 288 | |||
| Ablation 2: unified sequential | 328 | |||
| Ablation 3: w/o color compensation | 265 | |||
| Unified soft pruning (ours full) | 27.54 | 0.824 | 0.209 | 265 |
4.4 Ablation Studies
In this section, we conduct a series of ablation studies to validate the effectiveness of our proposed designs and answer two key questions:
1.
Is the replacement of spherical harmonics with spherical Gaussians necessary? And why not use the theoretically lighter-weight orthogonal-axis spherical variants? (Section 4.4.1)
2.
Is our proposed unified soft pruning framework effective? And does it offer a significant advantage over simply combining existing pruning methods? (Section 4.4.2)
4.4.1 Ablation on Color Modeling
To directly evaluate the effectiveness of our color modeling, we replace spherical harmonics (SH) in vanilla 3DGS with our arbitrarily-oriented spherical Gaussians (SG). We ensure a fair comparison by using identical hyperparameters and training strategies. As shown in Figure 3, our SG-based model achieves either higher quality with similar VRAM usage or comparable quality with significantly less VRAM compared to SH-based approaches.
Furthermore, we compare our arbitrarily-oriented SG to the orthogonal-axis SG proposed in SG-Splatting (Wang et al., 2024b). Although theoretically more lightweight, the fixed-axis variant suffers a significant drop in rendering quality (about 0.6 dB PSNR) as it cannot efficiently capture complex view-dependent effects. This also makes it unsuitable for lobe-pruning, as view-dependent effects cannot be guaranteed to align with its predefined axes.
4.4.2 Ablation on Pruning Strategies
To evaluate the efficacy of our unified soft pruning framework, we compare our final method with key ablation variants. First, we examine a version that ablates our lobe soft pruning, denoted as ”Ablation 1: soft hard” , which only performs constrained optimization on opacity, rather than on both opacity and sharpness. As shown in Table 4, the lack of a sharpness adjustment stage leads to an irreversible drop in performance when lobes are removed.
Next, we evaluate a variant that performs primitive and lobe pruning independently (sequential pruning), denoted as ”Ablation 2: unified sequential”. This sequential approach is inferior to our final unified method, which achieves a better rendering quality and VRAM trade-off. As we discussed in Section 3.3.2, this underscores the fundamental advantage of our unified framework. We have also experimentally validated the effectiveness of the introduced color compensation mechanism.
Finally, we compare our framework to two simple baselines to demonstrate that a naive combination of existing techniques is insufficient to achieve our goal. The first, which combines the SOTA pruning method GaussianSpa (Zhang et al., 2025b) with the spherical harmonics coefficient pruning of Reduced3DGS (Papantonakis et al., 2024), shows a severe rendering quality degradation (Table 4). This highlights the failure of existing spherical harmonics pruning metrics when applied to a reduced number of Gaussian primitives, which further validates our discussion on the limitations of spherical harmonics in Section 1. The second baseline, which replaces the spherical harmonics in GaussianSpa with spherical Gaussians, performs worse than our method in both memory footprint and rendering quality. This demonstrates that our pruning strategy not only reduces memory but also acts as a regularizer, improving rendering quality on novel views.
5 Limitation
Our current compression framework primarily focuses on static memory consumption to ensure generalizability across various rendering platforms and implementations. The dynamic portion of VRAM, which is highly dependent on specific renderer configurations (e.g., the tile-depth-Gaussian key-value table in a tile-based CUDA renderer), was not integrated into our framework as this remains a promising direction for future research. Furthermore, the ability of our framework to accurately model highlights with particularly complex or intricate shapes warrants more comprehensive testing and discussion.
6 Conclusion
In this paper, we introduced MEGS2, a novel framework that fundamentally solves the memory bottleneck of 3D Gaussian Splatting by jointly optimizing both primitive count and per-primitive parameters. Our approach pioneers a lightweight spherical Gaussian representation for color and a unified soft pruning method, which together deliver unprecedented memory compression. Our framework establishes a new paradigm for 3DGS compression by prioritizing rendering VRAM efficiency, a critical step towards enabling high-quality neural rendering on the edge.

### v2

@@ A.3 More Quantitive Results
| Scene | Method | PSNR | SSIM | LPIPS | Static VRAM | Rendering VRAM |
| Bonsai | 3DGS | 32.23 | 0.940 | 0.205 | 393 | 798 |
| Reduced3DGS | 31.44 | 0.933 | 0.214 | 66 | 276 | |
| GaussianSpa | 31.76 | 0.943 | 0.198 | 146 | 420 | |
| Ours | 31.95 | 0.943 | 0.192 | 44 | 266 | |
| 31.51 | 0.937 | 0.205 | 28 | 227 | ||
| Counter | 3DGS | 29.09 | 0.906 | 0.201 | 393 | 893 |
| Reduced3DGS | 28.57 | 0.899 | 0.212 | 67 | 331 | |
| GaussianSpa | 28.98 | 0.911 | 0.191 | 175 | 518 | |
| Ours | 28.78 | 0.906 | 0.195 | 44 | 326 | |
| 28.35 | 0.892 | 0.220 | 25 | 272 | ||
| Kitchen | 3DGS | 31.24 | 0.925 | 0.126 | 581 | 1197 |
| Reduced3DGS | 31.03 | 0.921 | 0.133 | 118 | 428 | |
| GaussianSpa | 31.50 | 0.929 | 0.127 | 163 | 443 | |
| Ours | 31.51 | 0.927 | 0.127 | 51 | 309 | |
| 31.27 | 0.923 | 0.133 | 36 | 275 | ||
| Room | 3DGS | 31.55 | 0.918 | 0.220 | 474 | 1007 |
| Reduced3DGS | 31.03 | 0.914 | 0.227 | 68 | 310 | |
| GaussianSpa | 31.61 | 0.922 | 0.211 | 141 | 401 | |
| Ours | 31.56 | 0.923 | 0.207 | 36 | 250 | |
| 31.16 | 0.915 | 0.226 | 22 | 214 |
| Scene | Method | PSNR | SSIM | LPIPS | Static VRAM | Rendering VRAM |
| Train | 3DGS | 21.97 | 0.818 | 0.198 | 397 | 773 |
| Reduced3DGS | 21.74 | 0.804 | 0.219 | 74 | 227 | |
| GaussianSpa | 21.81 | 0.826 | 0.197 | 164 | 339 | |
| Ours | 21.52 | 0.820 | 0.197 | 49 | 208 | |
| 21.32 | 0.818 | 0.205 | 35 | 160 | ||
| Truck | 3DGS | 25.39 | 0.881 | 0.143 | 752 | 1268 |
| Reduced3DGS | 25.28 | 0.875 | 0.154 | 105 | 276 | |
| GaussianSpa | 25.65 | 0.887 | 0.126 | 115 | 448 | |
| Ours | 25.37 | 0.886 | 0.121 | 54 | 214 | |
| 25.23 | 0.883 | 0.130 | 39 | 167 |
| Scene | Method | PSNR | SSIM | LPIPS | Static VRAM | Rendering VRAM |
| Bicycle | 3DGS | 25.20 | 0.764 | 0.210 | 1794 | 2955 |
| Reduced3DGS | 25.10 | 0.763 | 0.219 | 314 | 686 | |
| GaussianSpa | 25.43 | 0.779 | 0.222 | 240 | 491 | |
| Ours | 25.34 | 0.782 | 0.206 | 63 | 248 | |
| 25.16 | 0.771 | 0.226 | 48 | 205 | ||
| Stump | 3DGS | 26.61 | 0.769 | 0.217 | 1585 | 2487 |
| Reduced3DGS | 26.75 | 0.778 | 0.218 | 318 | 663 | |
| GaussianSpa | 27.07 | 0.802 | 0.207 | 222 | 460 | |
| Ours | 27.19 | 0.803 | 0.198 | 60 | 233 | |
| 27.04 | 0.797 | 0.213 | 47 | 198 | ||
| Garden | 3DGS | 27.36 | 0.863 | 0.108 | 1494 | 2449 |
| Reduced3DGS | 27.14 | 0.858 | 0.116 | 303 | 684 | |
| GaussianSpa | 27.22 | 0.854 | 0.140 | 227 | 460 | |
| Ours | 27.17 | 0.858 | 0.127 | 78 | 273 | |
| 26.62 | 0.846 | 0.143 | 63 | 240 | ||
| Flowers | 3DGS | 21.49 | 0.602 | 0.338 | 1066 | 1716 |
| Reduced3DGS | 21.37 | 0.598 | 0.345 | 221 | 497 | |
| GaussianSpa | 21.60 | 0.624 | 0.332 | 209 | 451 | |
| Ours | 21.65 | 0.625 | 0.331 | 66 | 255 | |
| 21.37 | 0.609 | 0.344 | 48 | 203 | ||
| Treehill | 3DGS | 22.58 | 0.633 | 0.328 | 1192 | 1947 |
| Reduced3DGS | 22.41 | 0.630 | 0.337 | 241 | 563 | |
| GaussianSpa | 22.84 | 0.655 | 0.312 | 226 | 471 | |
| Ours | 22.74 | 0.649 | 0.313 | 59 | 233 | |
| 22.44 | 0.640 | 0.335 | 43 | 188 |
| Scene | Method | PSNR | SSIM | LPIPS | Static VRAM | Rendering VRAM |
| Playroom | 3DGS | 30.00 | 0.903 | 0.244 | 678 | 1197 |
| Reduced3DGS | 29.96 | 0.904 | 0.248 | 90 | 276 | |
| GaussianSpa | 30.48 | 0.914 | 0.239 | 137 | 320 | |
| Ours | 30.77 | 0.914 | 0.229 | 52 | 218 | |
| 30.58 | 0.912 | 0.243 | 30 | 167 | ||
| DrJohnson | 3DGS | 29.41 | 0.902 | 0.239 | 1129 | 1941 |
| Reduced3DGS | 29.24 | 0.901 | 0.247 | 152 | 404 | |
| GaussianSpa | 29.50 | 0.910 | 0.239 | 184 | 424 | |
| Ours | 29.56 | 0.909 | 0.237 | 56 | 269 | |
| 29.45 | 0.904 | 0.250 | 37 | 219 |
| Method | RTX3060 | Dimensity 9400+ | Snapdragon 8+ Gen 1 | Snapdragon 888 |
| 3DGS | 26.3 | 6.6 | cannot render | connot render |
| GaussianSpa | 165.0 | 31.4 | render error | render error |
| Ours | 165.0 | 91.0 | 120.9 | 60.1 |
A.4 More Qualitative results
A.5 The Use of Large Language Models (LLMs)
The article was refined using Gemini 2.5 Pro / Gemini 2.5 Flash / DeepSeek R1 to improve phrasing for natural English expression and correct grammatical errors.